where in the oscillation is this maximum reached?

Suppose a 200- g mass(0.20kg)is oscillated at the end of the spring up on a frictionless horizontal surface. The Spring can be stretched and compressed and has a Spring constant of 240N/m. It was originally stretched a distance of 12cm from its equilibrium. Where in this oscillation this maximum reached?
Answer #1

The initial potential energy is

PEmax = ( 1/2) k x2   = ( 1/2 ) ( 240 N / m ) ( 0.12 m )2 = 1.728 J

Maximum velocity of the mass is,

v = ( ) (x) = ( ) (0.12 m)   = 4.156 m/s

the mass will have maximum potential energy when it is at maximum amplitude. (farthest point from equilibrium)

and it will have maximum velocity when the object reaches equilibrium position.